# A General Class of Model-Free Dense Precision Matrix Estimators

### v1

@@ Appendix A ONLINE APPENDIX
$\max_{1\leq j\leq p}\|\tilde{\Theta}_{j}^{\prime}-\Theta_{j}^{\prime}\|_{2}$$\leq$ $\max_{1\leq j\leq p}\left|\frac{1}{\tilde{\tau}_{j}^{2}}-\frac{1}{\tau_{j}^{2}}\right|+\max_{1\leq j\leq p}\|\tilde{\alpha}_{j}-\alpha_{j}^{*}\|_{2}\max_{1\leq j\leq p}[\frac{1}{\tilde{\tau}_{j}^{2}}]$ (OA.89)
@@
by Theorem 2(ii)(iii) under Assumptions 1, 5. Then by (OA.53) and Assumption 5(iii)
$\max_{1\leq j\leq p}\|\alpha_{j}^{*}\|_{2}=O(\frac{\sqrt{K}}{\sqrt{\bar{\xi}}})=o(1).$ (OA.90)
So the third term on right side of (OA.89) converges to zero faster in probability than the first right side term in (OA.89) given Theorem 2. Then Remark 4 of Theorem 2 shows that the first term on the right side of (OA.89) converges to zero slower in probability than the second term on the same right side. So the main rate comes from the first term, which is the main diagonal term rate as shown in Theorem 2(i). Q.E.D
We now provide three corollaries for specific subcases of our general estimator. Note that we provide the specific case of adaptive PCR estimator of $\tau_{j}^{2}$, which is defined as $\hat{\tau}_{j,PCR}^{2}$. In that respect the rates $r_{w1},r_{w2}$ simplify considerably. In adaptive PCR case these rates are with wpa1
$r_{w1,PCR}=\max\left(\sqrt{\frac{logn+K}{n\delta_{n}}},\sqrt{\frac{K}{\bar{\xi}\delta_{n}}}\right),$ (OA.91)
and
$r_{w2,PCR}=\max\left(\sqrt{\frac{logn+K}{n\delta_{n}}},\sqrt{\frac{K}{\bar{\xi}}}\right).$ (OA.92)
See that rates (OA.91)(OA.92) can be obtained by seeing $\bar{r}\leq K,\bar{\Psi}/\bar{\eta}\leq 1,\bar{\eta}\geq\delta_{n}$ wpa1 as shown in Corollary 1. Corollary OA.7 below follows through replacing $r_{w1},r_{w2}$ in Theorem 2 with $r_{w1,PCR},r_{w2,PCR}$ respectively.
Corollary OA.7.
(i). Under Assumptions 1-5(i) and using $r_{w1,PCR},r_{w2,PCR}$ instead of $r_{w1},r_{w2}$ respectively, and $r_{n}/\delta_{n}\to 1$,
@@
$+$ $O_{p}\left(d_{2n}r_{w2,PCR}\right)+O_{p}\left(max(\frac{K^{3/2}}{n^{1/2}},K\frac{\max(\sqrt{logp},\sqrt{logn})}{n^{1/2}})\right).$
(ii). Under Assumptions 1-5, $r_{n}/\delta_{n}\to 1$,
@@
Here we consider the second subcase $p-1<n$ scenario of RRE estimator of $\tau_{j}^{2}$. We define it as $\hat{\tau}_{RRE,j}^{2}$. We define the following rate which will be used instead of $r_{w1},r_{w2}$, wpa1, which can be deduced from Corollary 2 proof,
$r_{RRE-l}:=\max(\sqrt{\frac{logn+p}{n}},\sqrt{\frac{K}{\bar{\xi}}}).$ (OA.93)
The proof is a simple application of Theorem 2.
Corollary OA.8.
(i).Under Assumptions 1-5(i) and using $r_{RRE-l}$ instead of both $r_{w1},r_{w2}$, setting $p/n=a_{2n}\to 0$, and $r_{n}/\delta_{n}\to 1$
@@
We now consider the third subcase of RRE estimator when $p-1>n$. Note that this subcase will be different than the others, and we use an estimator of $\tau_{j}^{2}$ which is different due to interpolation property of RRE in this case. First let us show the interpolation and how it affects the estimator $\tilde{\tau}_{j}^{2}$ proposed, and the need to modify. In case of $p-1>n$, RRE estimator becomes the $l_{2}$ norm interpolator. Define the minimum norm interpolator as in (1) of Bunea et al. (2022), for each $j=1,\cdots,p$ $\hat{\alpha}_{j}:=argmin\{\|\alpha_{j}\|_{2}:\|Y_{-j}\alpha_{j}-y_{j}\|_{2}=\min_{\gamma_{j}}\|Y_{-j}\gamma_{j}-y_{j}\|_{2}.\}$, Lemma 28 of Bunea et al. (2022) shows $\hat{\alpha}_{j}=Y_{-j}^{+}y_{j}=\hat{\alpha}_{RRE,j},$ where $Y_{-j}^{+}:p-1\times n$ is the Moore-Penrose inverse of $Y_{-j}$. Then when $p-1>n$ $Y_{-j}\hat{\alpha}_{j}-y_{j}=Y_{-j}Y_{-j}^{+}y_{j}-y_{j}=y_{j}-y_{j}=0,$ with $Y_{-j}Y_{-j}^{+}=I_{n}$ by (136) of Bunea et al. (2022). This is interpolation, and we cannot use the formula for $\tilde{\tau}_{j}^{2}$ since with interpolation it can be shown, $\tilde{\tau}_{j}^{2}=0$ in RRE in (12), with $p-1>n$ case. So we propose the following estimator, which can be considered a subcase of general $\tilde{\tau}_{j}^{2}$. Define
$\hat{\tau}_{INT,j}^{2}:=\frac{y_{j}^{\prime}y_{j}}{n}.$ (OA.94)
Corollary OA.9.
@@
Proof of Corollary OA.9. (i). By (OB.11)(OA.94) and triangle inequality
@@
Next, Step 1 in the proof of Theorem 2 with (A.16), and (A.28) provides the result.
(ii). Given Assumptions in part (i) with $\sqrt{K}d_{2n}/\sqrt{\bar{\xi}}\to 0$ provides the consistency.
(iii). This is the same as in the proof of Theorem 2 (iii). Q.E.D.
ONLINE APPENDIX B
In this part, we provide a precision matrix formula. Different versions with time series, and for second moments of the regressor matrix can also be found in Callot et al. (2021), Caner and Kock (2018). Its relation to a hidden factor model context is established here for the first time in the literature and in detail. We follow matrix algebra and establish a formula for each row of the precision matrix. This relation will be essential to build up our estimator. Define all assets in the portfolio as $y_{t}=(y_{jt},Y_{-jt}^{\prime})^{\prime}:p\times 1$. In case of $j=3$, this represents $y_{3t}$ in the third cell in the vector $y_{t}$ and $Y_{-3t}$ represents elements $1,2,4,\cdots,p$ in $y_{t}$.
@@
$\Theta_{j,j}:=(\Sigma_{j,j}-\Sigma_{j,-j}\Sigma_{-j,-j}^{-1}\Sigma_{-j,j})^{-1}.$ (OB.1)
@@
$\Theta_{j,-j}:=-\Theta_{j,j}\Sigma_{j,-j}\Sigma_{-j,-j}^{-1}.$ (OB.2)
@@
$\alpha_{j}^{*}=\Sigma_{-j,-j}^{+}\Sigma_{-j,j},$ (OB.3)
where $\Sigma_{-j,-j}^{+}$ is the Moore-Penrose inverse of $\Sigma_{-j,-j}$. We will simplify that expression. First,
$\Sigma_{-j,-j}=E[(Y_{-jt}-\mu_{-j})(Y_{-jt}-\mu_{-j})^{\prime}]=A_{-j}\Sigma_{f}A_{-j}^{\prime}+\Sigma_{U,-j},$ (OB.4)
@@
by Assumption 2-3. So (OB.3) simplifies to
$\alpha_{j}^{*}=\Sigma_{-j,-j}^{-1}\Sigma_{-j,j}.$ (OB.5)
@@
where the first equality uses $\eta_{jt}$ definition, and the second equality uses definitions of moment matrices, and the third one uses (OB.5). Given (OB.6) we can write the following infeasible linear regression model,
$y_{jt}-\mu_{j}=(Y_{-jt}-\mu_{-j})^{\prime}\alpha_{j}^{*}+\eta_{jt}.$ (OB.7)
Now we relate (OB.2) using (OB.5). Then
$\Theta_{j,-j}=-\Theta_{j,j}(\alpha_{j}^{*})^{\prime},$ (OB.8)
and then define $\tau_{j}^{2}:=E\eta_{jt}^{2}$. See that by (OB.6)(OB.7)
@@
$\tau_{j}^{2}$ $=$ $\Sigma_{j,j}-(\alpha_{j}^{*})^{\prime}\Sigma_{-j,-j}\alpha_{j}^{*}=\Sigma_{j,j}-\Sigma_{j,-j}\Sigma_{-j,-j}^{-1}\Sigma_{-j,j}$ (OB.9)
$=$ $1/\Theta_{j,j},$ (OB.10)
where we use (OB.5) for the second equality and (OB.1) for the third one. Clearly, also by (OB.5)(OB.9)
$\tau_{j}^{2}=\Sigma_{j,j}-\Sigma_{j,-j}\alpha_{j}^{*}.$ (OB.11)
We can form the $j$ th row of precision matrix by using $j$ the main diagonal term, via (OB.10)(OB.11)
$\Theta_{j,j}=\frac{1}{\Sigma_{j,j}-\Sigma_{j,-j}\alpha_{j}^{*}},$ (OB.12)
and the rest of that row, via (OB.8)(OB.10)
$\Theta_{j,-j}=\frac{-(\alpha_{j}^{*})^{\prime}}{\tau_{j}^{2}}.$ (OB.13)
We start simplifying $\alpha_{j}^{*}$ formula in case of hidden factors as in (2)(3). The proof for (i) here is also a detailed version of the proof of Lemma 7 and (57) of Bunea et al. (2022). This is repeated to show that our scenario fits into their framework. Note that $\Sigma_{f}:=E(f_{t}-Ef_{t})(f_{t}-Ef_{t})^{\prime}$. In that respect in (OB.5), by(2)-(3)
$\Sigma_{-j,j}$ $=$ $E[(Y_{-jt}-\mu_{-j})(y_{jt}-\mu_{j})]=E\{[A_{-j}(f_{t}-Ef_{t})+U_{-jt}][(f_{t}-Ef_{t})^{\prime}a_{j}+u_{jt}]\}$ (OB.14)
$=$ $A_{-j}\Sigma_{f}a_{j}+A_{-j}E[(f_{t}-Ef_{t})u_{jt}]+E[U_{-jt}(f_{t}-Ef_{t})^{\prime}]a_{j}+E[U_{-jt}u_{jt}]$
@@
by Assumption 1(iv). Then the same analysis in (OB.14) provides with $\bar{A}_{-j}:=A_{-j}\Sigma_{f}^{1/2}$,
$\Sigma_{-j,-j}=E[(Y_{-jt}-\mu_{-j})(Y_{-jt}-\mu_{-j})^{\prime}]=A_{-j}\Sigma_{f}A_{-j}^{\prime}+\Sigma_{U,-j}=\bar{A}_{-j}\bar{A}_{-j}^{{}^{\prime}}+\Sigma_{U,-j}.$ (OB.15)
@@
$\Sigma_{-j,-j}^{-1}$ $=$ $(A_{-j}\Sigma_{f}A_{-j}^{\prime}+\Sigma_{U,-j})^{-1}$ (OB.16)
@@
$=$ $\Sigma_{U,-j}^{-1}-\Sigma_{U,-j}^{-1}\bar{A}_{-j}(I_{K}+\bar{A}_{-j}^{\prime}\Sigma_{U,-j}^{-1}\bar{A}_{-j})^{-1}\bar{A}_{-j}^{\prime}\Sigma_{U,-j}^{-1},$
@@
$\bar{A}_{-j}^{+}\bar{A}_{-j}=\bar{A}_{-j}^{\prime}(\bar{A}_{-j}^{+})^{\prime}=I_{K},$ (OB.17)
since $A_{-j},\Sigma_{f}$ has full rank K by Assumption 3, so (135) of Bunea et al. (2022) provides the last result (OB.17). Define $\bar{a}_{j}:=\Sigma_{f}^{1/2}a_{j}$ and (OB.5)(OB.14)
$\alpha_{j}^{*}$ $=$ $\Sigma_{-j,-j}^{-1}\bar{A}_{-j}\bar{a}_{j}=\Sigma_{-j,-j}^{-1}[\bar{A}_{-j}\{\bar{A}_{-j}^{\prime}(\bar{A}_{-j}^{+})^{\prime}\}\bar{a}_{j}]$ (OB.18)
@@
where we use (OB.17) for the second equality and for the third equality (OB.15) and $\bar{A}_{-j}$ definition. By using $\bar{G}_{j}$ definition before Lemma 1 and using (OB.16) on right side of (OB.18) for $\Sigma_{-j,-j}^{-1}$
$I_{p-1}-\Sigma_{-j,-j}^{-1}\Sigma_{U,-j}$ $=$ $I_{p-1}-[\Sigma_{U,-j}^{-1}-\Sigma_{U,-j}^{-1}\bar{A}_{-j}\bar{G}_{j}^{-1}\bar{A}_{-j}^{\prime}\Sigma_{U,-j}^{-1}]\Sigma_{U,-j}$ (OB.19)
@@
Use (OB.19) in (OB.18)
$\alpha_{j}^{*}=\Sigma_{U,-j}^{-1}\bar{A}_{-j}\bar{G}_{j}^{-1}\bar{A}_{-j}^{\prime}(\bar{A}_{-j}^{+})^{\prime}\bar{a}_{j}=\Sigma_{U,-j}^{-1}\bar{A}_{-j}\bar{G}_{j}^{-1}\bar{a}_{j},$ (OB.20)
by (OB.17) for the second equality. One of the other elements in the precision matrix is the reciprocal of the main diagonal term, $\tau_{j}^{2}$ in (OB.11). We provide a formula for that in terms of hidden factor regression model, which will be new. In that respect, by (2)(3), Assumption 1
$\Sigma_{j,j}:=E(y_{jt}-\mu_{j})^{2}=a_{j}^{\prime}\Sigma_{f}a_{j}+\sigma_{j}^{2}=\bar{a}_{j}^{\prime}\bar{a}_{j}+\sigma_{j}^{2}.$ (OB.21)
@@
$\Sigma_{j,-j}:=E(y_{jt}-\mu_{j})(Y_{-jt}-\mu_{-j})^{\prime}=\bar{a}_{j}^{\prime}\bar{A}_{-j}^{\prime}.$ (OB.22)
Then combine (OB.20)(OB.21)(OB.22) in (OB.11)
$\tau_{j}^{2}=(\bar{a}_{j}^{\prime}\bar{a}_{j}+\sigma_{j}^{2})-\bar{a}_{j}^{\prime}\bar{A}_{-j}^{\prime}\Sigma_{U,-j}^{-1}\bar{A}_{-j}\bar{G}_{j}^{-1}\bar{a}_{j}.$ (OB.23)
(i)-(ii). By (OB.20) and (OB.23), and using them in (OB.10)(OB.13) we have the desired result. Q.E.D.
@@
where $tr(H)$ represents trace of $H$, and $\gamma_{e}$ subgaussian random vector is defined in Assumption 4. Now, we provide a lemma for a lower bound inequality for singular value of a specific random matrix that we use in our proofs for RRE when $p-1>n$. In that way we extend the zero factor mean in Proposition 14 of Bunea et al. (2022) to a nonzero mean case below.
Lemma OB.2.
Under Assumptions $3^{*}(i)$, Assumption 3(ii), Assumption 4, with some $C>0$ large enough, $re(\Sigma_{U,-j})>Cn$, for each $j=1,\cdots,p$, and $\tilde{U}_{-j}$ has independent entries, in the case of $p-1>n$
@@
Proof of Lemma OB.2. So first we use (89) in Bunea et al. (2022) and then equation below that on p.42 of proof of Proposition 14 of Bunea et al. (2022) but we replace Theorem 4.6.1 of Vershynin (2019) in that proof with Theorem 5.3.9 of Vershynin (2012), (set $t=\sqrt{n}$) to take into account nonzero mean of factors. The rest carries over from the proof of Proposition 14 of Bunea et al. (2022).Q.E.D.
ONLINE APPENDIX C
@@
By (2) and (3) and Assumption 4, for each $t=1,...,n$, we have:
$y_{t}=Af_{t}+u_{t},$ (OC.1)
@@
$[\Sigma_{f}]_{i,j}=(-1)^{i+j}(0.3)^{|i-j|}\min([\Sigma_{f}]_{i,i},[\Sigma_{f}]_{j,j}).$ (OC.2)
Each element of $A$ is drawn independently from the normal distribution $\mathcal{N}(0.5,1/\sqrt{K})$. The idiosyncratic error, $u_{t}$, is drawn from the normal distribution, $\mathcal{N}(0,\Sigma_{u})$.
In our design, we define $\Sigma_{u}:p\times p$ as a diagonal matrix with diagonal elements sampled from uniform distribution with $\text{Unif}(1,3)\times\log(p)$. With that we capture noisy setups in complex models and can understand the effect of noise on adaptive PCR and RRE with $p-1>n$ specifically. The results for $K=3,5$ (true number of hidden factors) are presented in Tables OC.1-OC.2, respectively. We also have the same design with true number of latent factors as $K=20$, the results were very similar to $K=5$ hence not reported. This $K=20$ Table can be obtained from authors on demand.However note that we still use $K=20$ in other tables to see the effects of fitting larger number of factors when the true DGP has few factors. In addition, to further demonstrate how our methods behave as the signal-to-noise ratio (SNR) changes, we rewrite (OC.1) in compact form:
$Y=FA+U$ (OC.3)
Similar to Bing et al. (2021), we multiply each element of $A$ by a scalar $\alpha$ chosen from the range ${0.1,...,1.5}$. Fixing $K=20,n=400$, we experiment in high- dimensional case, with $p=450$. The low dimensional case is almost identical to high dimensional case, so we report only high dimensional case. Then for each $\alpha$, we calculate the SNR and plot the precision matrix estimation error for any of the utilized models. The results are presented in Figure 1.

### v2

@@ 6 Conclusion
$\max_{1\leq j\leq p}\|\tilde{\Theta}_{j}^{\prime}-\Theta_{j}^{\prime}\|_{2}$ $\leq$ $\max_{1\leq j\leq p}\left|\frac{1}{\tilde{\tau}_{j}^{2}}-\frac{1}{\tau_{j}^{2}}\right|+\max_{1\leq j\leq p}\|\tilde{\alpha}_{j}-\alpha_{j}^{*}\|_{2}\max_{1\leq j\leq p}[\frac{1}{\tilde{\tau}_{j}^{2}}]$ (A.116)
@@
by Theorem 2(ii)(iii) under Assumptions 1, 5. Then by (A.53) and Assumption 5(iii)
$\max_{1\leq j\leq p}\|\alpha_{j}^{*}\|_{2}=O(\frac{\sqrt{K}}{\sqrt{\bar{\xi}}})=o(1).$ (A.117)
So the third term on right side of (A.116) converges to zero faster in probability than the first right side term and the first term converges to zero faster than the second term in (A.116) given Theorem 2. Q.E.D
We now provide three corollaries for specific subcases of our general estimator. Note that we provide the specific case of adaptive PCR estimator of $\tau_{j}^{2}$, which is defined as $\hat{\tau}_{j,PCR}^{2}$. In that respect the rate $r_{w1}$ simplifies considerably. In adaptive PCR case these rates are with wpa1
$r_{w1,PCR}=\max\left(\sqrt{\frac{logn+K}{n}},\sqrt{\frac{K}{\bar{\xi}}}\right),$ (A.118)
See that rate (A.118) can be obtained by seeing $\bar{r}\leq K,\bar{\Psi}/\bar{\eta}\leq 1,\bar{\eta}\geq\beta$ wpa1 as shown in Corollary 1. Corollary A.1 below follows through replacing $r_{w1}$ in Theorem 2 with $r_{w1,PCR}$.
Corollary A.1
(i). Under Assumptions 1-5(i) and using $r_{w1,PCR}$ instead of $r_{w1}$,
@@
$+$ $O_{p}\left(d_{2n}r_{w1,PCR}\right)+O_{p}\left(max(\frac{K^{3/2}}{n^{1/2}},K\frac{\max(\sqrt{logp},\sqrt{logn})}{n^{1/2}})\right).$
(ii). Under Assumptions 1-5,
@@
Here we consider the second subcase p1<np-1<n scenario of RRE estimator of $\tau_{j}^{2}$. We define it as $\hat{\tau}_{RRE,j}^{2}$. We define the following rate which will be used instead of $r_{w1}$, wpa1, which can be deduced from Corollary 2 proof,
$r_{RRE-l}:=\max(\sqrt{\frac{logn+p}{n}},\sqrt{\frac{K}{\bar{\xi}}}).$ (A.119)
Corollary A.2
(i).Under Assumptions 1-5(i) and using $r_{RRE-l}$ instead of both $r_{w1}$, setting $p/n=a_{2n}\to 0$,
@@
We now consider the third subcase of RRE estimator when p1>np-1>n. Note that this subcase will be different than the others, and we use an estimator of $\tau_{j}^{2}$ which is different due to interpolation property of RRE in this case. First let us show the interpolation and how it affects the estimator $\tilde{\tau}_{j}^{2}$ proposed, and the need to modify. In case of p1>np-1>n, RRE estimator becomes the $l_{2}$ norm interpolator. Define the minimum norm interpolator as in (1) of Bunea et al. (2022), for each $j=1,\cdots,p$
$\hat{\alpha}_{j}:=argmin\{\|\alpha_{j}\|_{2}:\|Y_{-j}\alpha_{j}-y_{j}\|_{2}=\min_{\gamma_{j}}\|Y_{-j}\gamma_{j}-y_{j}\|_{2}.\},$
Lemma 28 of Bunea et al. (2022) shows $\hat{\alpha}_{j}=Y_{-j}^{+}y_{j}=\hat{\alpha}_{RRE,j}.$ Then when p1>np-1>n,
$Y_{-j}\hat{\alpha}_{j}-y_{j}=Y_{-j}Y_{-j}^{+}y_{j}-y_{j}=y_{j}-y_{j}=0,$
with $Y_{-j}Y_{-j}^{+}=I_{n}$ by (136) of Bunea et al. (2022). This is interpolation, and we cannot use the formula for $\tilde{\tau}_{j}^{2}$ since with interpolation it can be shown, $\tilde{\tau}_{j}^{2}=0$ in RRE in (11), with p1>np-1>n case. So we propose the following estimator, which can be considered a subcase of general $\tilde{\tau}_{j}^{2}$. Define
$\hat{\tau}_{INT,j}^{2}:=\frac{y_{j}^{\prime}y_{j}}{n}.$ (A.120)
Corollary A.3
@@
Proof of Corollary A.3. (i). By (B.11)(A.120) and triangle inequality
@@
Next, Step 1 in the proof of Theorem 2 with (A.16), and (A.28) provides the result. (ii). Given Assumptions in part (i) with $\sqrt{K}d_{2n}/\sqrt{\bar{\xi}}\to 0$ provides the consistency. (iii). This is the same as in the proof of Theorem 2 (iii). Q.E.D.
APPENDIX B
In this part B, we provide a precision matrix formula and then provide two lemmata on tail inequalities.
Precision Matrix Formula Derivation
Precision matrix relation to a hidden factor model context is established here for the first time in the literature and in detail. Define all assets in the portfolio as $y_{t}=(y_{jt},Y_{-jt}^{\prime})^{\prime}:p\times 1$. In case of $j=3$, this represents $y_{3t}$ in the third cell in the vector $y_{t}$ and $Y_{-3t}$ represents elements $1,2,4,\cdots,p$ in $y_{t}$.
@@
$\Theta_{j,j}:=(\Sigma_{j,j}-\Sigma_{j,-j}\Sigma_{-j,-j}^{-1}\Sigma_{-j,j})^{-1}.$ (B.1)
@@
$\Theta_{j,-j}:=-\Theta_{j,j}\Sigma_{j,-j}\Sigma_{-j,-j}^{-1}.$ (B.2)
@@
$\alpha_{j}^{*}=\Sigma_{-j,-j}^{+}\Sigma_{-j,j},$ (B.3)
where $\Sigma_{-j,-j}^{+}$ is the Moore-Penrose inverse of $\Sigma_{-j,-j}$. First,
$\Sigma_{-j,-j}=E[(Y_{-jt}-\mu_{-j})(Y_{-jt}-\mu_{-j})^{\prime}]=A_{-j}\Sigma_{f}A_{-j}^{\prime}+\Sigma_{U,-j},$ (B.4)
@@
by Assumption 2-3. (B.3) simplifies to
$\alpha_{j}^{*}=\Sigma_{-j,-j}^{-1}\Sigma_{-j,j}.$ (B.5)
@@
where the first equality uses $\eta_{jt}$ definition, and the second equality uses definitions of moment matrices, and the third one uses (B.5). Given (B.6) we can write the following infeasible linear regression model,
$y_{jt}-\mu_{j}=(Y_{-jt}-\mu_{-j})^{\prime}\alpha_{j}^{*}+\eta_{jt}.$ (B.7)
Now we relate (B.2) using (B.5). Then
$\Theta_{j,-j}=-\Theta_{j,j}(\alpha_{j}^{*})^{\prime},$ (B.8)
and then define $\tau_{j}^{2}:=E\eta_{jt}^{2}$. See that by (B.6)(B.7)
@@
$\tau_{j}^{2}$ $=$ $\Sigma_{j,j}-(\alpha_{j}^{*})^{\prime}\Sigma_{-j,-j}\alpha_{j}^{*}=\Sigma_{j,j}-\Sigma_{j,-j}\Sigma_{-j,-j}^{-1}\Sigma_{-j,j}$ (B.9)
$=$ $1/\Theta_{j,j},$ (B.10)
where we use (B.5) for the second equality and (B.1) for the third one. Clearly, also by (B.5)(B.9)
$\tau_{j}^{2}=\Sigma_{j,j}-\Sigma_{j,-j}\alpha_{j}^{*}.$ (B.11)
We can form the $j$ th row of precision matrix by using $j$ the main diagonal term, via (B.10)(B.11)
$\Theta_{j,j}=\frac{1}{\Sigma_{j,j}-\Sigma_{j,-j}\alpha_{j}^{*}},$ (B.12)
and the rest of that row, via (B.8)(B.10)
$\Theta_{j,-j}=\frac{-(\alpha_{j}^{*})^{\prime}}{\tau_{j}^{2}}.$ (B.13)
We start simplifying $\alpha_{j}^{*}$ formula in case of hidden factors as in (2)(3). The proof for (i) here is also a detailed version of the proof of Lemma 7 and (57) of Bunea et al. (2022). This is repeated to show that our scenario fits into their framework. Note that $\Sigma_{f}:=E(f_{t}-Ef_{t})(f_{t}-Ef_{t})^{\prime}$. In that respect in (B.5), by(2)-(3)
$\Sigma_{-j,j}=E[(Y_{-jt}-\mu_{-j})(y_{jt}-\mu_{j})]$ $=$ $E\{[A_{-j}(f_{t}-Ef_{t})+U_{-jt}][(f_{t}-Ef_{t})^{\prime}a_{j}+u_{jt}]\}$ (B.14)
$=$ $A_{-j}\Sigma_{f}a_{j}+A_{-j}E[(f_{t}-Ef_{t})u_{jt}]$
$+$ $E[U_{-jt}(f_{t}-Ef_{t})^{\prime}]a_{j}+E[U_{-jt}u_{jt}]$
@@
by Assumption 1(iv). Then the same analysis in (B.14) provides with $\bar{A}_{-j}:=A_{-j}\Sigma_{f}^{1/2}$,
$\Sigma_{-j,-j}=E[(Y_{-jt}-\mu_{-j})(Y_{-jt}-\mu_{-j})^{\prime}]=A_{-j}\Sigma_{f}A_{-j}^{\prime}+\Sigma_{U,-j}=\bar{A}_{-j}\bar{A}_{-j}^{{}^{\prime}}+\Sigma_{U,-j}.$ (B.15)
@@
$\Sigma_{-j,-j}^{-1}$ $=$ $(A_{-j}\Sigma_{f}A_{-j}^{\prime}+\Sigma_{U,-j})^{-1}$ (B.16)
@@
$=$ $=\Sigma_{U,-j}^{-1}-\Sigma_{U,-j}^{-1}\bar{A}_{-j}(I_{K}+\bar{A}_{-j}^{\prime}\Sigma_{U,-j}^{-1}\bar{A}_{-j})^{-1}\bar{A}_{-j}^{\prime}\Sigma_{U,-j}^{-1},$
@@
$\bar{A}_{-j}^{+}\bar{A}_{-j}=\bar{A}_{-j}^{\prime}(\bar{A}_{-j}^{+})^{\prime}=I_{K},$ (B.17)
since $A_{-j},\Sigma_{f}$ has full rank K by Assumption 3, so (135) of Bunea et al. (2022) provides the last result (B.17). Define $\bar{a}_{j}:=\Sigma_{f}^{1/2}a_{j}$ and (B.5)(B.14)
$\alpha_{j}^{*}$ $=$ $\Sigma_{-j,-j}^{-1}\bar{A}_{-j}\bar{a}_{j}=\Sigma_{-j,-j}^{-1}[\bar{A}_{-j}\{\bar{A}_{-j}^{\prime}(\bar{A}_{-j}^{+})^{\prime}\}\bar{a}_{j}]$ (B.18)
@@
where we use (B.17) for the second equality and for the third equality (B.15) and $\bar{A}_{-j}$ definition. By using $\bar{G}_{j}$ definition before Lemma 1 and using (B.16) on right side of (B.18) for $\Sigma_{-j,-j}^{-1}$
$I_{p-1}-\Sigma_{-j,-j}^{-1}\Sigma_{U,-j}$ $=$ $I_{p-1}-[\Sigma_{U,-j}^{-1}-\Sigma_{U,-j}^{-1}\bar{A}_{-j}\bar{G}_{j}^{-1}\bar{A}_{-j}^{\prime}\Sigma_{U,-j}^{-1}]\Sigma_{U,-j}$ (B.19)
@@
Use (B.19) in (B.18)
$\alpha_{j}^{*}=\Sigma_{U,-j}^{-1}\bar{A}_{-j}\bar{G}_{j}^{-1}\bar{A}_{-j}^{\prime}(\bar{A}_{-j}^{+})^{\prime}\bar{a}_{j}=\Sigma_{U,-j}^{-1}\bar{A}_{-j}\bar{G}_{j}^{-1}\bar{a}_{j},$ (B.20)
by (B.17) for the second equality. One of the other elements in the precision matrix is the reciprocal of the main diagonal term, $\tau_{j}^{2}$ in (B.11). We provide a formula for that in terms of hidden factor regression model, which will be new. In that respect, by (2)(3), Assumption 1
$\Sigma_{j,j}:=E(y_{jt}-\mu_{j})^{2}=a_{j}^{\prime}\Sigma_{f}a_{j}+\sigma_{j}^{2}=\bar{a}_{j}^{\prime}\bar{a}_{j}+\sigma_{j}^{2}.$ (B.21)
@@
$\Sigma_{j,-j}:=E(y_{jt}-\mu_{j})(Y_{-jt}-\mu_{-j})^{\prime}=\bar{a}_{j}^{\prime}\bar{A}_{-j}^{\prime}.$ (B.22)
Then combine (B.20)(B.21)(B.22) in (B.11)
$\tau_{j}^{2}=(\bar{a}_{j}^{\prime}\bar{a}_{j}+\sigma_{j}^{2})-\bar{a}_{j}^{\prime}\bar{A}_{-j}^{\prime}\Sigma_{U,-j}^{-1}\bar{A}_{-j}\bar{G}_{j}^{-1}\bar{a}_{j}.$ (B.23)
(i)-(ii). By (B.20) and (B.23), and using them in (B.10)(B.13) we have the desired result. Q.E.D.
Tail Inequalities
@@
where $tr(H)$ represents trace of $H$, and $\gamma_{e}$ subgaussian random vector is defined in Assumption 4. Now, we provide a lemma for a lower bound inequality for singular value of a specific random matrix for RRE when p1>np-1>n. This extends the zero factor mean in Proposition 14 of Bunea et al. (2022) to a nonzero mean case below.
Lemma B.2
Under Assumptions $3^{*}(i)$, Assumption 3(ii), Assumption 4, with some $C>0$ large enough, $re(\Sigma_{U,-j})>Cn$, for each $j=1,\cdots,p$, and $\tilde{U}_{-j}$ has independent entries, in the case of p1>np-1>n
@@
Proof of Lemma B.2. So first we use (89) in Bunea et al. (2022) and then equation below that on p.42 of proof of Proposition 14 of Bunea et al. (2022) but we replace Theorem 4.6.1 of Vershynin (2019) in that proof with Theorem 5.3.9 of Vershynin (2012), (set $t=\sqrt{n}$) to take into account nonzero mean of factors. The rest carries over from the proof of Proposition 14 of Bunea et al. (2022).Q.E.D.
APPENDIX C
In this part C, we cover simulations and extra empirics without transaction costs and discussion of literature and hidden factors.
@@
By (2) and (3) and Assumption 4, for each t=1,,nt=1,...,n, we have:
$y_{t}=Af_{t}+u_{t},$
@@
$\Sigma_{f}]_{i,j}=(-1)^{i+j}(0.3)^{|i-j|}\min([\Sigma_{f}]_{i,i},[\Sigma_{f}]_{j,j}).$
Each element of $A$ is drawn independently from the normal distribution $\mathcal{N}(0.5,1/\sqrt{K})$. The idiosyncratic error, $u_{t}$, is drawn from the normal distribution, $\mathcal{N}(0,\Sigma_{u})$. In our design, we define $\Sigma_{u}:p\times p$ as a diagonal matrix with diagonal elements sampled from uniform distribution with $\text{Unif}(1,3)\times\log(p)$. With that we capture noisy setups in complex models and can understand the effect of noise on adaptive PCR and RRE with p1>np-1>n specifically. The results for $K=3,5$ (true number of hidden factors) are presented in Table C.1. However note that we also use $K=20$ to see the effects of fitting larger number of factors when the true DGP has few factors.